KCH 16 Traveling Waves Problem 16.35 Review A concert loudspeaker suspended high off the ground emits 29.0 W of sound power. A small microphone with a 1.00 cm area is 55.0 m from the speaker Part A What is the sound intensity at the position of the microphone? Express your answer with the appropriate units. Value Units Submit Part B How much sound energy impinges on the microphone each second? Express your answer with the appropriate units. LA Value Units Submit Provide Feedback
Answer #1

A)

Intensity, I = P/ 4 pi r^2

I = 29/(4 pi x 55^2) = 7.63 x 10^-4 W/m^2

B)

Energy, E = I a = (7.63 x 10^-4)(1 x 10^-4) = 7.63 x 10^-8 J/s

Comment in case any doubt please rate my answer.. ..